OpenStudy (anonymous):

A colony of bacteria grows according to the uninhibited growth model. Suppose there is 18 g of bacteria on Monday and 54 g of bacteria on Wednesday. (a) Find a function that gives the amount of bacteria in grams after t days. (Use the variable a to represent the initial amount.) A(t) = (b) What is the doubling time for the colony? (Round your answer to 2 decimal places.) days A colony of bacteria grows according to the uninhibited growth model. Suppose there is 18 g of bacteria on Monday and 54 g of bacteria on Wednesday. (a) Find a function that gives the amount of bacteria in grams after t days. (Use the variable a to represent the initial amount.) A(t) = (b) What is the doubling time for the colony? (Round your answer to 2 decimal places.) days @Mathematics

OpenStudy (anonymous):

\[\frac{54}{18}=3\] so it triples every three days. function will be \[A(t)=18\times 3^{\frac{t}{3}}\]

OpenStudy (anonymous):

if you want to know the doubling time set \[2=3^{\frac{t}{3}}\] and solve for t via \[\frac{t}{3}=\frac{\ln(2)}{\ln(3)}\] or \[t=\frac{3\ln(2)}{\ln(3)}\]

OpenStudy (anonymous):

first answer aint working